# mystyle backgroundcolor=mygray, leftmargin=0.45em, rightmargin=0.45em, innertopmargin=10pt, innerbottommargin=10pt, skipabove=10pt, skipbelow=10pt

### v3

@@ IV-D RQ4: How does the key code extraction impact detection performance?
Experimental Results. We present the experimental results of utilizing Key Code Extraction on SeVC in Table IX and DiverseVul in Figure 6. We find that there is a noticeable improvement for SeVC (i.e., 98.5% vs 92.4% for Single-Type and 83.8% vs 64.4% for Multi-Type). For DiverseVul, there is a back-and-forth trend comparing the results with key code and without key code. The potential reason might be that the longer and more complex code length makes it difficult to effectively extract the key code on DiverseVul. The results indicate that this enhancement scheme can considerably improve the detection accuracy of the model in different vulnerability explanation tasks. This demonstrates the importance of supplementing code semantic information for the fine-tuning of large vulnerability models and the feasibility of semantic information extraction based on large models. Key code extraction offers two major benefits: (1) Enhanced Focus: By isolating the most relevant code segments, the model can focus on critical information, leading to more accurate detection results. (2) Noise Reduction: Eliminating irrelevant code improves the signal-to-noise ratio, enabling the model to better learn and detect vulnerability patterns. Moreover, this approach relies on automated annotation rather than manual feature extraction, making it broadly applicable and scalable. Our experimental results preliminarily validate the feasibility and utility of key code extraction through LLMs.
Figure 6: Performance of Key Code on DiverseVul(RQ4).
Summary of RQ4: By guiding LLMs to focus on key code, the performance of vulnerability detection can be significantly improved(e.g., from 64.4% to 83.8% for Multi-Type detection on SeVC).
@@ V Discussion
Implication 1: Broader Applications for Vulnerability-Related Tasks. Our study reveals that integrating vulnerability explanation with detection does not compromise the model’s performance. This finding suggests that we should consider a wider range of vulnerability-related tasks beyond mere detection when fine-tuning LLMs. By doing so, we can potentially enhance the model’s overall understanding and its ability to provide more contextually rich and actionable insights. For instance, tasks such as vulnerability impact assessment, prioritization based on severity, or even automated patch generation could be integrated into the training regime. This holistic approach could lead to the development of more sophisticated tools that not only identify vulnerabilities but also assist in managing and mitigating associated risks.
Implication 2: Effectiveness of LLMVulExp Using Different LLMs. In this paper, we use CodeLlama-13B-Instruct as the primary model to conduct the experiments. To investigate the effectiveness of our framework using different LLMs, we utilize two additional LLMs, CodeLlama-7B-Instruct and Llama3-8B-Instruct, to conduct the experiments following the setting of RQ1.3. Figure 7 shows the results of multi-type vulnerability detection using LLMs of different sizes. We find that LLMVulExp can achieve effective performance with these additional LLMs. Notably, the performance of CodeLlama-13B-Instruct and CodeLlama-7B-Instruct is similar and outperforms Llama3-8B-Instruct. Overall, LLMVulExp demonstrates its effectiveness using different LLMs of various sizes.
Figure 7: Performance of different base LLMs.
Implication 3: Dependency on Annotation Quality in Fine-Tuning Frameworks. The effectiveness of fine-tuning heavily depends on the quality of the annotations used for training. Our study highlights the critical role that high-quality annotations play in the fine-tuning process. It is essential to develop robust annotation frameworks that ensure consistency and accuracy. Furthermore, the annotation process itself could be enhanced through the use of semi-automated tools that provide initial labels, which are then reviewed and refined by human experts. This hybrid approach could balance the need for detailed annotations and the practical constraints of manual labeling.
@@ VI Threats to Validity
Construct Validity. We evaluate explanatory capability using three effective metrics, but these metrics may not cover all aspects of explanations. To mitigate bias, we use both manual and automated reviews with LLMs. Two authors independently annotate the explanations, resolving discrepancies until consensus is reached, achieving a Cohen’s Kappa [46] value of 0.76, indicating substantial agreement.
Internal Validity. Due to the high computational cost of fine-tuning and inference evaluation, we couldn’t conduct experiments under different data splits and randomization states. Although we limit the number of hyperparameter trials for LoRa configuration and didn’t fine-tune larger models like the 34B version due to GPU constraints, these limitations don’t undermine our objective: exploring the potential of LLMs for vulnerability detection and explanation.
External Validity. We choose the advanced open-source code LLM CodeLlama and evaluate different versions as well as the recently released Llama3 on two distinct datasets. These datasets, while containing a significant amount of vulnerable code, include only a few CWE types and mostly generic vulnerabilities, which do not fully reflect the diversity and complexity of vulnerabilities found in real-world development environments. Although these datasets contain substantial vulnerable code, they differ from real development environments, posing a risk of limited generalization.
@@ VII Conclusion
In this paper, we propose LLMVulExp, a framework designed to detect and explain vulnerabilities using LLMs. The results underscore the potential of LLMs in advancing vulnerability detection and explanation in software security. Our research provides valuable insights into the fine-tuning of LLMs for vulnerability detection and explanation. It highlights the importance of addressing the data volume bottleneck for training vulnerability LLMs in software development. Furthermore, the quality of annotations is paramount for the success of fine-tuning frameworks. Considering these insights, future work can aim to develop more capable and efficient models that significantly contribute to the field of software security.

### v4

@@ IV-D RQ4: How does the key code extraction impact detection and explanation performance?
TABLE XII: Impact of Key Code on Detection Performance on the PrimeVul Dataset (RQ4).
Metric Without Key Code With Key Code
Multi-Class CWE Detection
Precision (Macro) 19.2 19.3
Recall (Macro) 25.7 27.6
F1 Score (Macro) 20.7 21.5
Binary Vulnerability Detection
Precision 29.5 31.7
Recall 61.1 62.6
F1 Score 39.8 42.1
Figure 6: Impact of Key Code on Detection Performance on the DiverseVul Dataset (RQ4).
Experimental Results. Our findings indicate that the effectiveness of key code extraction is highly dependent on the complexity of the dataset.
On SeVC—the dataset featuring synthesized data and relatively simple vulnerability classification—key code extraction yields notably superior performance as demonstrated in Table XI. On the more challenging PrimeVul dataset, key code provides a clear advantage. As shown in Table XII, the “With Key Code” model consistently outperforms the baseline. The F1 score for binary detection improves from 39.8% to 42.1%, and the Macro-F1 for multi-class detection increases from 20.7% to 21.5%. This positive result is significant because PrimeVul is the benchmark where models struggled most (RQ1.4). The improvement stems from two benefits: (1) Enhanced Focus, which directs the model’s attention to the most relevant code, and (2) Noise Reduction, which improves the signal-to-noise ratio by filtering out irrelevant context. This demonstrates that for complex, noisy, real-world code, guiding the model’s focus is a highly effective strategy.
On the DiverseVul dataset, the impact is inconclusive. In contrast, the results illustrated in Figure 6 show a mixed trend, with neither model establishing a definitive advantage. A plausible explanation is that the vulnerabilities in DiverseVul are more straightforward, allowing the baseline model to perform well without explicit guidance. In such cases, the overhead of the key code extraction process offers no significant additional benefit.
Regarding explanation quality, our manual and LLM evaluation investigation that simply providing key code during fine-tuning did not yield a significant improvement in the quality of generated explanations. This suggests that a more sophisticated integration strategy is needed to translate focused attention into higher-quality explanatory text. Exploring how to better leverage key code to enhance explanation quality remains an important direction for future research.
Summary of RQ4: Key code extraction enhances vulnerability detection performance on challenging real-world datasets by effectively guiding the model’s attention mechanisms.
@@ V Discussion
Implication 1: Broader Applications for Vulnerability-Related Tasks. Our finding that integrating vulnerability explanation with detection preserves model performance suggests expanding LLM fine-tuning beyond detection. Incorporating tasks like impact assessment, severity prioritization, or automated patch generation could enhance contextual understanding and actionability. This holistic approach may yield tools that simultaneously identify and mitigate vulnerabilities.
Implication 2: Effectiveness of LLMVulExp Using Different LLMs. In this paper, we use CodeLlama-13B-Instruct as the primary model to conduct the experiments. A key question is how our framework’s effectiveness varies with different base LLMs. We investigate this from both smaller and larger directions.
First, we utilize two additional widely-used smaller LLMs, CodeLlama-7B-Instruct and Llama3-8B-Instruct, on the easier DiverseVul dataset. As shown in Figure 7, LLMVulExp achieves effective and comparable performance with these models, demonstrating that our proposed fine-tuning and explanation generation strategies are not limited to a single model architecture and can be successfully applied to smaller LLMs. Second, we seek to understand if simply employing a larger, more advanced model could further boost performance, particularly on a more challenging, realistic dataset. To this end, we conduct experiments on the PrimeVul dataset, comparing the performance of CodeLlama-13B-Instruct against a significantly larger and more powerful model, DeepSeek-Coder-33B-Instruct. The results, presented in Figure 8, lead to a crucial insight: the more advanced model did not yield a discernible accuracy gain. This finding consistently suggests that for fine-tuning tasks on current, small-scale vulnerability datasets, simply increasing the base model’s size or using a more recent architecture may not be the silver bullet. We hypothesize that the primary bottleneck for improving performance is now the scale and diversity of the training data, rather than the inherent capacity of the base model. This is a very valuable insight for the field, suggesting that future efforts should focus on data-centric approaches to unlock the full potential of large models for vulnerability detection.
Figure 7: Different base LLMs on DiverseVul.
Figure 8: Different base LLMs on PrimeVul.
Implication 3: Dependency on Annotation Quality in Fine-Tuning Frameworks. To demonstrate the high sensitivity of fine-tuning effectiveness to annotation quality, we conduct experiments using different annotation LLMs. Specifically, CodeLlama-7B fine-tuned with DeepSeek-V3-generated explanations achieved an average Macro-F1 of 20.5% on PrimeVul, while the same model fine-tuned with CodeLlama-13B-generated explanations attained only 16.4%. This significant performance gap substantiates the critical role of annotation quality. This dependency underscores the necessity of developing robust annotation frameworks that ensure consistency and accuracy. A promising approach involves semi-automated tools where advanced LLMs provide initial annotations subsequently refined by human experts. Such a hybrid methodology could mitigate quality variations while balancing the practical constraints of manual curation.
@@ VI Threats to Validity
Construct Validity. We evaluated explanatory capability using three metrics (Accuracy, Clarity, and Actionability). While informative, these metrics may not be exhaustive, and further refinement is desirable. To mitigate bias, two authors independently annotated explanations, achieving substantial agreement. Furthermore, our use of GPT-3.5 and DeepSeek-V3 for certain automated annotation tasks introduces a replication challenge: its specific version may deprecate or evolve, potentially impacting result reproducibility. Future work addressing this should meticulously document model versions and prompts or consider open-source alternatives. This overall rigorous approach supports our evaluation’s reliability, though we acknowledge this specific consideration for LLM-assisted annotation components.
Internal Validity. Due to the high computational cost of fine-tuning and inference evaluation of LLMs, we are unable to conduct experiments under different data splits and randomization states. The randomness in our experiments (e.g., data splitting, explanation annotation, and fine-tuning process) may affect the results. However, we conduct multiple trials to validate the stability of our conclusions for each experimental setup. Although the number of hyperparameter trials for LoRa configuration and the LLM generation method are relatively limited, and due to GPU constraints, we do not attempt to fine-tune larger models, we ensure that these limitations did not fundamentally undermine our primary objective: to explore the potential of conducting vulnerability detection and explanation tasks using LLMs.
External Validity. We chose the advanced open-source code LLM CodeLlama, based on the Llama2 architecture, as our primary model for this research. We test different versions, including 7B and 13B, and evaluate the performance of newer LLM Llama3 and DeepSeek-Coder on vulnerability detection and explanation tasks. However, there are many different architectures of general-purpose open-source LLMs and code models, and our experiments are limited to a few models. Additionally, we conduct experiments using three differently configured open-source project-based vulnerability datasets. Although these datasets include a relatively large amount of vulnerable code, there are still differences compared to code in real development environments. Consequently, our experimental conclusions have an inherent risk of not generalizing to other vulnerability datasets or industrial environments. Nonetheless, our careful selection and thorough evaluation process aim to ensure that our findings remain relevant and insightful within the scope of our study.
@@ VII Conclusion
In this paper, we propose LLMVulExp, a framework designed to detect and explain vulnerabilities using LLMs. The results underscore the potential of LLMs in advancing vulnerability detection and explanation in software security.
Our research provides valuable insights into the fine-tuning of LLMs for vulnerability detection and explanation. It highlights the importance of addressing the data volume bottleneck for training vulnerability LLMs in software development. Furthermore, the quality of annotations is paramount for the success of fine-tuning frameworks. Considering these insights, future work can aim to develop more capable and efficient models that significantly contribute to the field of software security.
Acknowledgment
This work was supported by the National Key R&D Program of China (No. 2024YFB4506400).